OpenStudy (anonymous):

Write the ratio of the diameter of ball A to the diameter of ball B diameter of ball A: 44mm diameter of ball B: 7.2 cm the diameter written as a simplified fraction

OpenStudy (sleepyhead314):

"the ratio of the diameter of ball A to the diameter of ball B" (diameter of ball A) -> ----------------- = ? (diameter of ball B)

OpenStudy (anonymous):

Ratio will be: \[\frac{Diameter \; of \; A}{Diameter \; of \; B} = \frac{44 \; mm}{7.2 \; cm}\]

OpenStudy (anonymous):

its all different units.. must get to the same unit or she will take points off

OpenStudy (anonymous):

It must be noted that : you should make the units of the two diameters equal first..

OpenStudy (sleepyhead314):

oh right like 1 cm = 10mm

OpenStudy (anonymous):

\[mm = 10^{-3}m\] \[cm = 10^{-2} m\]

OpenStudy (anonymous):

exactly i went towards cm.. put it back into a fraction and got 11/18

OpenStudy (anonymous):

exactly i went towards cm.. put it back into a fraction and got 11/18

OpenStudy (sleepyhead314):

that ratio is correct :)

OpenStudy (anonymous):

Ratio will be: \[\frac{4.4}{7.2} \implies \frac{11}{18}\]

OpenStudy (anonymous):

Good.. :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You are welcome dear..